English
Tamil Nadu Board of Secondary EducationHSC Science Class 12

Two balls are drawn in succession without replacement from an urn containing four red balls and three black balls. Let X be the possible outcomes drawing red balls. Find the probability mass

Advertisements
Advertisements

Question

Two balls are drawn in succession without replacement from an urn containing four red balls and three black balls. Let X be the possible outcomes drawing red balls. Find the probability mass function and mean for X

Chart
Sum
Advertisements

Solution

Let X be the random variable that denotes the number of red balls.

X = {0, 1, 2}

Sample space consist of 7C2 = 21

X= 0, X-1 (BB) = 3C2

X= 1, X-1 (BR) = 3C1 × 4C1 = 12

X = 2, X-1 (RR) 4C2 = 6

Values of random variable 0 1 2 Total
Number of elements in inverse image 3 12 6 21

Probability mass function

x 0 1 2
F(x) `3/21` `12/21` `6/21`

Mean: `mu = "E"("X")`

= `sum x f(x)`

= `0 xx 3/21 + 1 xx 12/21 + 2 xx 6/21`

= `24/21`

= `8/7`

shaalaa.com
Mathematical Expectation
  Is there an error in this question or solution?
Chapter 11: Probability Distributions - Exercise 11.4 [Page 210]

APPEARS IN

Samacheer Kalvi Mathematics - Volume 1 and 2 [English] Class 12 TN Board
Chapter 11 Probability Distributions
Exercise 11.4 | Q 2 | Page 210

RELATED QUESTIONS

For the random variable X with the given probability mass function as below, find the mean and variance.

`f(x) = {{:((4 - x)/6,  x = 1","  2","  3),(0,  "otherwise"):}`


For the random variable X with the given probability mass function as below, find the mean and variance.

`f(x) = {{:(2(x - 1), 1 < x ≤ 2),(0, "otherwise"):}`


The time to failure in thousands of hours of an electronic equipment used in a manufactured computer has the density function
`f(x) = {{:(3"e"^(-3x), x > 0),(0, "eleswhere"):}`
Find the expected life of this electronic equipment


A lottery with 600 tickets gives one prize of ₹ 200, four prizes of ₹ 100, and six prizes of ₹ 50. If the ticket costs is ₹ 2, find the expected winning amount of a ticket


Choose the correct alternative:

Consider a game where the player tosses a six-sided fair die. If the face that comes up is 6, the player wins ₹ 36, otherwise he loses ₹ k2, where k is the face that comes up k = {1, 2, 3, 4, 5}. The expected amount to win at this game in ₹ is


Choose the correct alternative:

On a multiple-choice exam with 3 possible destructive for each of the 5 questions, the probability that a student will get 4 or more correct answers just by guessing is


Choose the correct alternative:

If P(X = 0) = 1 – P(X = 1). If E[X] = 3 Var(X), then P(X = 0) is


Choose the correct alternative:

If X is a binomial random variable with I expected value 6 and variance 2.4, Then P(X = 5) is


Let X be a continuous random variable with probability density function
`"f"_x(x) = {{:(2x",", 0 ≤ x ≤ 1),(0",",  "otherwise"):}`
Find the expected value of X


Let X be a continuous random variable with probability density function
f(x) = `{{:(3/x^4",",  x ≥ 1),(0",",  "otherwise"):}`
Find the mean and variance of X


What are the properties of Mathematical expectation?


A person tosses a coin and is to receive ₹ 4 for a head and is to pay ₹ 2 for a tail. Find the expectation and variance of his gains


Choose the correct alternative:

Given E(X) = 5 and E(Y) = – 2, then E(X – Y) is


Choose the correct alternative:

E[X – E(X)]2 is


Choose the correct alternative: 

`int_(-oo)^oo` f(x) dx is always equal to


Choose the correct alternative: 

An expected value of a random variable is equal to it’s


Choose the correct alternative: 

The distribution function F(x) is equal to


Share
Notifications

Englishहिंदीमराठी


      Forgot password?
Use app×